2.4 A force F is applied at point A of the block with edge lengths of a, b, c Determine (1) the projections of F on the x, y and z axes; (2) the moments of F about x, y and z axes; (3) the moment of F about OB axis.
1
Expert's answer
2021-09-22T00:22:41-0400



(1)


"F_x=-|\vec F|\sin \beta\cos \alpha"

"F_y=-|\vec F|\cos \beta\cos \alpha"

"F_z=|\vec F|\sin \alpha"

(2)


"\vec F=(-|\vec F|\sin \beta\sin \alpha, -|\vec F|\sin \beta\cos \alpha, |\vec F|\cos \beta)"

"\vec r=(a, b, c)"

"\vec M_O=\vec r \times\vec F"

"=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
a & b & c\\
-|\vec F|\sin \beta\cos \alpha & -|\vec F|\cos \beta\cos \alpha & |\vec F|\sin \alpha\\
\end{vmatrix}"

"=( |\vec F|b\sin \alpha+|\vec F|c\cos \beta\cos \alpha)\vec i"

"+(- |\vec F|a\sin \alpha-|\vec F|c\sin \beta\cos \alpha)\vec j"

"+(- |\vec F|a\cos \alpha\cos \beta+|\vec F|b\sin \beta\cos \alpha)\vec k"

"M_x=\vec i\cdot(\vec r\times\vec F)="

"= |\vec F|b\sin \alpha+|\vec F|c\cos \beta\cos \alpha"

"M_y=\vec j\cdot(\vec r\times\vec F)="

"=- |\vec F|a\sin \alpha-|\vec F|c\sin \beta\cos \alpha"

"M_z=\vec k\cdot(\vec r\times\vec F)="

"- |\vec F|a\cos \alpha\cos \beta+|\vec F|b\sin \beta\cos \alpha"

(3)


"\overrightarrow{OB}=(a, 0, c))"

"\vec u=(\dfrac{a}{\sqrt{a^2+b^2}}, 0, \dfrac{c}{\sqrt{a^2+c^2}})"

"M_{OB}=\vec u\cdot(\vec r\times\vec F)="

"=\dfrac{ |\vec F|ab\sin \alpha+|\vec F|ac\cos \beta\cos \alpha}{\sqrt{a^2+b^2}}"

"+\dfrac{- |\vec F|ac\cos \alpha\cos \beta+|\vec F|bc\sin \beta\cos \alpha}{\sqrt{a^2+b^2}}"

"=\dfrac{ |\vec F|ab\sin \alpha+|\vec F|bc\sin \beta\cos \alpha}{\sqrt{a^2+b^2}}"